# Direct imaging and Talbot Phase Imaging at the Matter in Extreme Conditions X-ray Imager of LCLS

### v2

@@ V Talbot Phase Imaging (TPI)
which can be efficiently implement using a fast Fourier transform and its inverse.
We tested this method on a shock wave in a silicon sample driven by the MEC short pulse laser. First an image was taken of the illumination (i.e. the beam without a sample in place) to use as a a reference. Then a measurement of the shock in the sample was taken. Figure 10.a) shows the raw data of the shock, while b) shows the differential phase in the y direction. The reconstructed intensity is shown in Figure10.c). As can be seen, the structure of the Be lenses is highly visible in that image. In Figure10.d) we show the actual phase introduced by the sample; the structure in the intensity introduced by the lenses disappears, as it is not present in the phase. The phase change behind the elastic wave can clearly be seen, as well as the larger shift after the phase change.
The Fourier analysis of the initial 2d fringe pattern limits the resolution that can be attained due Nyquist–Shannon sampling theorem. In the data shown in Figure 10 the resolution of the phase measurement to approximately 1.2m. However, the imaging geometry at MEC has the capability of magnification of up to 40x at photon energies below 10keV. Since we can easily resolve a 4um talbot pitch on the scintilator based detector, a resolution of 8um in the phase measurement at the detector or nm is easily achievable.
@@ VI Concluding remarks
In conclusion, we have developed at X-ray imaging diagnostic at MEC that is available to the user community. It can be used in combination with the MEC long pulse laser to, for example, image shock waves and phase transformation in materials, and with the MEC short pulse laser system, to image its interaction with dense matter. The diagnostic has a 200nm resolution, and can operate using photon energies ranging from 6keV to 24keV. In combination with with Talbot grating, a Talbot Imaging setup can be used that is capable of measuring both intensity and phase changes introduced by a target with 200nm resolution.
@@ Acknowledgements.
A.E.G. acknowledges support from the LANL Reines LDRD and from the Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division under contract DE-AC02-76SF00515 (FWP 100320). Use of the Linac Coherent Light Source (LCLS), SLAC National Accelerator Laboratory, is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515. The MEC instrument and B.N., E.G., D.K., G.D. and H.J.L are supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences under contract DE-AC02-76SF00515.References

### v4

@@ V Talbot Imaging
which can be efficiently implemented using a fast Fourier transform and its inverse.
We used this method on a shock compressed silicon sample driven by the MEC long pulse laser, as detailed in Figure 10.
First an image was taken of the white field illumination (i.e., the beam without a sample in place) to use as a reference. Then a measurement of the shock in the sample was taken. Figure 10 10(a) shows the raw data of the shock, while 10 10(b) shows the differential phase in the y direction. The reconstructed intensity is shown in Figure 10 10(c). The structure of the Be lenses is highly visible in that image. In Figure 10 10(d) we show the actual phase introduced by the sample; the structure in the intensity introduced by the lenses disappears, as it is not present in the phase. The phase change behind the elastic wave can clearly be seen, as well as the larger shift after the phase change.
The Fourier analysis of the initial two dimensional fringe pattern limits the resolution that can be attained due to the Nyquist–Shannon sampling theorem. In the data shown in Figure 10 the resolution of the phase measurement is approximately 1.2 m. However, the imaging geometry at MEC has the capability of magnification of up to 40 at photon energies below 10 keV. Since we can easily resolve a 4 m Talbot pitch on the scintillator based detector, a resolution of 8 m in the phase measurement at the detector or nm is achievable.
@@ VI Concluding remarks
In conclusion, we have developed an X-ray imaging diagnostic at MEC that is available to the user community. It can be used in combination with the MEC long pulse laser to, for example, image shock waves and phase transformation in materials over large field of view, and with the MEC short pulse laser system, to image its interaction with dense plasmas at high resolution. The diagnostic has a 200 nm spatial resolution, taking 50 fs (the pulse duration of the FEL) snapshot of excited HED samples, and can operate using photon energies ranging from 5 to 24 keV. In combination with the direct imaging geometry, a Talbot Imaging setup can be used to measure both intensity and phase changes introduced by a target with sub-micron resolution.
@@ Acknowledgements.
A.E.G. acknowledges support from the LANL Reines LDRD and from the DOE FES ECA, 2019. Use of the Linac Coherent Light Source (LCLS) SLAC National Accelerator Laboratory, was supported by the U.S. Department of Energy Office of Science, Basic Energy Science under Contract No. DE-AC02-76SF00515. The MEC instrument is supported also by the U.S. Department of Energy Office of Science, Fusion Energy Science under Contract No. DE-AC02-76SF00515, FWP 100106.References
@@
- Dewald et al. (2018) E. L. Dewald, O. L. Landen, L. Masse, D. Ho, Y. Ping, D. Thorn, N. Izumi, L. Berzak Hopkins, J. Kroll, A. Nikroo, and J. A. Koch, Review of Scientific Instruments 89, 10G108 (2018), https://pubs.aip.org/aip/rsi/article-pdf/doi/10.1063/1.5039346/16727284/10g108_1_online.pdf .
- Do et al. (2022) A. Do, C. R. Weber, E. L. Dewald, D. T. Casey, D. S. Clark, S. F. Khan, O. L. Landen, A. G. MacPhee, and V. A. Smalyuk, Phys. Rev. Lett. 129, 215003 (2022).
@@
- Oliver et al. (2022) M. Oliver, C. H. Allen, L. Divol, Z. Karmiol, O. L. Landen, Y. Ping, R. Wallace, M. Schölmerich, W. Theobald, T. Döppner, and T. G. White, Review of Scientific Instruments 93, 093502 (2022), https://pubs.aip.org/aip/rsi/article-pdf/doi/10.1063/5.0091348/16601249/093502_1_online.pdf .
- Morace et al. (2014) A. Morace, L. Fedeli, D. Batani, S. Baton, F. N. Beg, S. Hulin, L. C. Jarrott, A. Margarit, M. Nakai, M. Nakatsutsumi, P. Nicolai, N. Piovella, M. S. Wei, X. Vaisseau, L. Volpe, and J. J. Santos, Physics of Plasmas 21, 102712 (2014), https://pubs.aip.org/aip/pop/article-pdf/doi/10.1063/1.4900867/14866977/102712_1_online.pdf .
@@
- (28) Optique Peter, (2012), https://www.optiquepeter.com/en/index.php.
- (29) Advanced hCMOS Systems, (2021), https://www.hcmos.com/.
- Hart et al. (2019) P. A. Hart, A. Carpenter, L. Claus, D. Damiani, M. Dayton, F.-J. Decker, A. Gleason, P. Heimann, E. Hurd, E. McBride, S. Nelson, M. Sanchez, S. Song, and D. Zhu, in X-Ray Free-Electron Lasers: Advances in Source Development and Instrumentation V, Vol. 11038, edited by T. Tschentscher, L. Patthey, K. Tiedtke, and M. Zangrando, International Society for Optics and Photonics (SPIE, 2019) p. 110380Q.
- Looker et al. (2020) Q. Looker, A. P. Colombo, M. Kimmel, and J. L. Porter, Review of Scientific Instruments 91, 043502 (2020), https://pubs.aip.org/aip/rsi/article-pdf/doi/10.1063/5.0004711/14731052/043502_1_online.pdf .
- Hodge et al. (2022) D. S. Hodge, A. F. T. Leong, S. Pandolfi, K. Kurzer-Ogul, D. S. Montgomery, H. Aluie, C. Bolme, T. Carver, E. Cunningham, C. B. Curry, M. Dayton, F.-J. Decker, E. Galtier, P. Hart, D. Khaghani, H. J. Lee, K. Li, Y. Liu, K. Ramos, J. Shang, S. Vetter, B. Nagler, R. L. Sandberg, and A. E. Gleason, Opt. Express 30, 38405 (2022).
- Decker et al. (2022) F.-J. Decker, K. L. Bane, W. Colocho, S. Gilevich, A. Marinelli, J. C. Sheppard, J. L. Turner, J. J. Turner, S. L. Vetter, A. Halavanau, C. Pellegrini, and A. A. Lutman, Scientific Reports 12, 3253 (2022).
- Halavanau et al. (2021) A. Halavanau, S. Carbajo, F.-J. Decker, A. Krasnykh, A. Lutman, A. Marinelli, C. Mayes, and D. Nguyen, in Proc. IPAC’21, International Particle Accelerator Conference No. 12 (JACoW Publishing, Geneva, Switzerland, 2021) pp. 365–368.
@@
- Hagemann et al. (2021) J. Hagemann, M. Vassholz, H. Hoeppe, M. Osterhoff, J. M. Rosselló, R. Mettin, F. Seiboth, A. Schropp, J. Möller, J. Hallmann, C. Kim, M. Scholz, U. Boesenberg, R. Schaffer, A. Zozulya, W. Lu, R. Shayduk, A. Madsen, C. G. Schroer, and T. Salditt, Journal of Synchrotron Radiation 28, 52 (2021).
@@
- Nagler et al. (2017) B. Nagler, A. Aquila, S. Boutet, E. C. Galtier, A. Hashim, M. S. Hunter, M. Liang, A. E. Sakdinawat, C. G. Schroer, A. Schropp, M. H. Seaberg, F. Seiboth, T. van Driel, Z. Xing, Y. Liu, and H. J. Lee, Scientific Reports 7, 13698 (2017).
- Makita et al. (2020) M. Makita, G. Seniutinas, M. H. Seaberg, H. J. Lee, E. C. Galtier, M. Liang, A. Aquila, S. Boutet, A. Hashim, M. S. Hunter, T. van Driel, U. Zastrau, C. David, and B. Nagler, Optica 7, 404 (2020).